Lura:

Use the table to write a proportion. MONDAY TUESDAY GALLONS 6 8 MILES 180 m (It wouldn't let me post in math)

carmelle:

Set up your proportion as: \[\frac{ Gallons }{ Miles }\] \[\frac{ 6 }{ 180 } = \frac{ 8 }{ m }\] The next step is to cross multiply, can you do that?

Lura:

yes.thank you

carmelle:

Alright, can you do that or do you still need help?